A fast –food chain wishes to determine whether there are any differences in three media (magazine, TV, radio) in terms of consumer recall of an ad. The results of an advertising study are as follows:

a.At the 0.05 level of significance, determine whether there is evidence of a significant media effect with respect to the proportion of individuals who can recall the ad. b.If appropriate, use the .05 level of significance and determine which media are different in recall ability of the ad.

Recall Ability

Medium

Magazine

TV

Radio

Total

Remember Ad

25

10

7

42

Do Not Remember Ad

73

93

108

274

Total

98

103

115

316
Answer #1
Column 1 Column 2 Column 3 Total
Row 1 25 10 7 42
Row 2 73 93 108 274
Total 98 103 115 316
Expected Values Column 1 Column 2 Column 3 Total
Row 1 98*42/316 = 13.025 103*42/316 = 13.69 115*42/316 = 15.285 42
Row 2 98*274/316 = 84.975 103*274/316 = 89.31 115*274/316 = 99.715 274
Total 98 103 115 316

Null and Alternative Hypotheses

H0: there are no significance  differences in three media (magazine, TV, radio) in terms of consumer recall of an ad

Ha: there are significance  differences in three media (magazine, TV, radio) in terms of consumer recall of an ad

Rejection Region Based on the information provided, the significance level is a 0.05 , the number of degrees of freedom is df- (2-1) x (3-1) 2, so then the rejection region for this test is R : χ2 > 5.991) The Chi-Squared statistic is computed as follows: X (og-Eij). = 11.000+ 1.687 + 0.995 + 0.152 + 4.491 + 0.688 = 19.022 Decision about the null hvpothesis Since it is observed that χ. = 19.022 > χ。= 5.991, it is then concluded that the null hypothesis is rejected

Conclusion : At  0.05 level of significance there is evidence of a significant media effect with respect to the proportion of individuals who can recall the ad.